# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ III.3 Energy dressing
Returning to the exchange field, we observed that the FLT Rabi rate can be interpreted as the projection of the exchange field onto the axis perpendicular to the spin in the plane containing both the exchange field and the spin itself. However, we initially ignored the projection of this field along the spin axis. On the quantization axis of the spin the Hamiltonian is diagonal, meaning that the projection of the exchange on this axis will change the energy of the states, usually referred to as dressing of the energy. This projection provides a Zeeman-like splitting even in the absence of an external magnetic field, as long as the exchange field and spin are not perpendicular. When an external magnetic field is applied, the energy shift will modify the resonance frequency of the spectrum, shifting it relative to the bare resonance frequency given by $f_{0}=\Delta_{eg}/h=g\mu_{B}B/h$. Following this geometric interpretation, we define the energy (or frequency) shift as:
$\Delta E(t)$ $=$ $h\Delta f(t)=g\mu_{B}B_{\mathrm{exch}}(t)\cos\theta$
@@
The time-dependent component in Eq. LABEL:energy_dress does not affect the spin dynamics, as it aligns with the quantization axis, though we include it here for completeness.
@@
This relation was used to derive Eqs. LABEL:spin_equations. Equation 31 provides a straightforward method to effectively measure the Rabi frequency, given the knowledge of the frequency shift and the angle between the tip and the external magnetic field. Another important detail about the energy dressing is how it affects zero-harmonic and higher-harmonic signals. Writing Eq. 5 in Floquet components, as in [35, 36], shows that in general the resonance frequency per harmonic is $hf_{n\omega}=(\Delta_{eg}+h\Delta f)/n$ which makes the energy shift dependent on the harmonic (or Floquet number) $n$. Interestingly, if we move to the perspective of the magnetic field being swept at a fixed frequency, the resonance magnetic field per harmonic will be $B_{n\omega}=h(nf-\Delta f)/g\mu_{B}$. This implies that the shifting in the magnetic field domain is the same per harmonic since $\Delta f$ is not been altered by $n$. As a consequence, the zero-harmonic feature in the frequency domain appears at zero frequency, while in the magnetic field domain, it does at $-h\Delta f/g\mu_{B}=-B_{\mathrm{exch}}\cos\theta$. This offset can be used to measure the exchange field indirectly, a proposal in line with other studies [42].
@@ IV Results
The left lead represents the spin-polarized tip with a spin polarization of $P_{L}=0.8$ along with the findings in [30]. The coupling strengths are set to $\gamma_{R}^{0}=3\times\gamma_{L}^{0}=15\ \mu$eV for all simulations and the temperature is equal to $T=1$ K for both the left and right leads. In the context of ESR-STM experiments, this implies that the tip-sample distance is constant and the background current will change accordingly with the DC bias. The DC bias drop is applied to the tip so $eV_{DC}=\mu_{L}$ since we align the chemical potential of the right lead with the Fermi energy, which defines the zero energy. The singly occupied energy is $\varepsilon=-100$ meV, and the Coulomb repulsion is $U=250$ meV. These values offer a realistic enough exploration of both tunneling regimes not so far off from DFT and others predictions [25, 72]. We have increased the intrinsic width of the impurity states to $\hbar/\tau_{c}=0.25$meV, compared to the one used in [36], allowing exploration of the CB regime while maintaining $\hbar/\tau_{c}\ll|\varepsilon|,\ \varepsilon+U$. The driving amplitude is set to $A_{L}=50\%$, applied only to the polarized lead, consistent with the experimental setup. Finally, the Zeeman splitting provides a bare resonance frequency of approximately $f_{0}=16.8$ GHz, with an isotropic $g$-factor of 2. The energy dressing will modify the resonance position to a value $f_{0}^{\prime}=\Delta f+f_{0}$, see Eq. LABEL:energy_dress, which is taken into account to perform on resonance simulations. Additional parameters are specified in figure captions.
IV.1 Spin accumulation vs. DC bias without AC driving
@@
The effect of spin accumulation persists even in the CB regime, albeit weaker, vanishing completely at zero voltage. There is, however, a competition between the relaxation time, which ranges from 0.2 to 50 ns (Fig. 7a), and $\hbar/\dot{S}_{\mathrm{acc}}$, leading to different results depending on the DC bias. In the CB regime, $\hbar/\dot{S}_{\mathrm{acc}}\gg\tau_{\mathrm{rel}}$, so the spin evolution is dominated by the relaxation term, as the accumulation term acts extremely slowly on the QI. This opens the possibility of an FLT-ESR signal since the time dependent exchange field will have ”time” to torque the spin coherently, producing Rabi oscillations. On the other hand, above the energy thresholds outside of the CB regime, both the relaxation and spin accumulation terms are of the same order of magnitude, forcing the QI to be spin-accumulated in the quantization axis before any coherent evolution occurs. Therefore, AC driving that leads to a time-dependent spin accumulation will provide an incoherent STT-ESR signal, as we will see in the following sections.
@@ IV.2 ESR in the STT-driven regime
Figure 3: ESR outside of the CB regime; out of plane magnetic field. a) Bloch sphere representation of the spin components in the lab frame for $\theta=\pi/18$ and positive $eV_{\mathrm{DC}}/|\varepsilon|=2$. The spin is initialized thermally at $\langle\mathbf{S}\rangle_{\mathrm{thermal}}=(-0.033,0,-0.188)$ (red arrow). The positive DC bias causes the spin to flip to its spin polarization in the steady state as a consequence the spin accumulation term. At the same time, the spin continuously performs Larmor oscillations even in the steady state. The Bloch sphere axes are reduced for clarity. b) Time-evolution of the current from the thermal state, reaching the steady state after 2 ns. c) Corresponding ESR electric current signal in the STT-driven regime in the frequency domain. The resonance frequency is slightly different to the bare one at 16.8 GHz being now $\Delta f+f_{0}=17.07$ GHz due to the energy dressing.
Figure 4: ESR outside of the CB regime; in-plane magnetic field. a) Bloch sphere representation of the spin components in the lab frame for $\theta=\pi/2$ and $eV_{\mathrm{DC}}/|\varepsilon|=2$. The spin is thermally initialized at $\langle\mathbf{S}\rangle_{\mathrm{thermal}}=(-0.19,0,0)$ (red arrow). Here, the spin accumulation forces alignment with the quantization axis of the spin in the steady state, leading to continuous Larmor oscillations in the $zy$-plane. b) Time-evolution of the current from the thermal situation. c) ESR electric current signal in the STT regime for the case of a) in the frequency domain. Since in-plane initialization leads to not dressing, the resonance occurs at the bare one $f_{0}=16.8$ GHz.
@@
Fig. 3b illustrates the spin dynamics, characterized by fast Larmor oscillations that persist even in the long-time limit. These oscillations occur in a plane perpendicular to the quantization axis, Fig. 3a, defined by the magnetic field, and arise from the time-dependent spin accumulation. This accumulation, proportional to the polarized current, attempts to torque the accumulated spin toward the spin polarization axis, which in result makes the crossed product $\mathbf{S}\times\mathbf{B}$ non-zero, forcing the spin to torque with an intensity proportional to $\sin\theta$. In this regime, the populations remain frozen, and the Larmor oscillations only modify the coherences, making the ESR current fully governed by them [36].
@@
Fig. 3c shows the corresponding frequency behavior. Although the ESR signal is significant, the background current is three orders of magnitude higher, resulting in an extremely low ESR signal relative to the background. This is expected in the STT-driven regime, as reflected in the low-quality factor $\Omega_{\mathrm{STT}}T_{2}$. The signal exhibits an asymmetric Fano profile due to the exchange field and the residual presence of homodyne detection. However, it becomes fully symmetric when $eV_{\mathrm{DC}}\gg\varepsilon+U$, where the exchange field vanishes (see Eq. 25).
@@
For the in-plane configuration, Fig. 4a, the strongest Larmor oscillations and ESR signals (Figs. 4b and 4c)) are achieved. The Larmor precession now occurs in the $zy$-plane, again perpendicular to the magnetic field and the quantization axis. If the STT-driven regime would follow the homodyne detection previously reported [34, 21], the ESR signal would be minimal or zero since the spin and tip spin polarization are perpendicular, and the ESR signal is proportional to their scalar product. However, in this regime, this limitation does not hold as the spin will be driven towards the $z$ axis by the spin accumulation. After this initial push, the spin feels the torque from the external magnetic field, from the cross product $\langle\mathbf{S}\rangle\times\mathbf{B}$. This torque is maximized for in plane configurations, producing a stronger ESR signal compared to the out of plane situation. This interaction allows the initial oscillatory evolution of the spin observed both in Fig. 4a and Fig. 3a. Shortly thereafter, the spin settles into oscillations within the $zy$ plane in Fig. 4a, driven by the time-dependent spin accumulation and the torque exerted by the external magnetic field.
@@
The role of the exchange field in this scenario is analogous to the out-of-plane case; however, the ESR signal in Fig. 4c is nearly symmetrical. This symmetry arises because, when the spin is initialized along a Cartesian axis, $\langle S_{z}\rangle$ in Eq. 13 depends solely on either the populations or the decoherences, but not on a combination of both. Consequently, no Fano profile emerges. Specifically, for an in-plane magnetic field, the $z$ component of the spin depends only on the real part of the coherences and the $x$ one gives the populations. Importantly, with an initial in-plane configuration, the spin cannot polarize because the spin accumulation term is fully perpendicular to the external magnetic field. As a result, the system tends to equalize populations into a superposition state at the center of the Bloch sphere if no driving is applied. Note that in the in-plane configuration, the populations, measured in the $x$ axis, also remain unperturbed in the long-time limit, even at resonance, as observed in the out-of-plane situation Fig. 3.
@@
The reader may notice that the ESR signal in both cases (Figs. 3c and 4c) are dips (a reduction in the current on-resonance) which is a direct result of the STT-driven regime at positive bias. In this regime, the coherences peak on resonance while remaining zero off resonance. According to Eq. 12, this behavior produces a dip in the signal. Conversely, under a negative DC bias ($eV_{\mathrm{DC}}<\varepsilon$), the current becomes negative and the ESR signal will flip from a dip to a peak. However, the ESR signal cannot flip by merely changing the sign of the spin polarization because an additional sign change occurs in the $z$ component of the spin, effectively preserving the overall response. Higher order tunneling processes could induce additional flipping of the signal though.
@@ IV.3 ESR in the FLT-driven regime
We now shift our focus to the Coulomb blockade (CB) regime, where FLT-driven resonance dominates. This regime favors quantum-coherent manipulation of the spin over the spin polarization of the impurity spin since the spin accumulation term is significantly minimized. This reduces the impact of the completely decoherent polarized current on the impurity, enabling the spin to be used as a qubit. FLT driven regime is basically equivalent to conventional ESR where a time dependent magnetic field induces resonant transition. As highlighted in the theoretical discussion, this allows for quality factors $\Omega_{\mathrm{FLT}}T_{2}$ to reach their upper limits, constrained only by the ability to enhance $T_{2}$ for a fixed Rabi rate. Notably, this enhancement is strongly dependent on the DC bias, making it the critical factor in achieving high coherence for spin manipulation.
@@
Figure 5: ESR in the Coulomb Blockade regime; out of plane magnetic field. a) Bloch sphere representation of the spin components in the rotating frame for a magnetic field tilted by $\theta=\pi/18$ and a bias voltage of $eV_{\mathrm{DC}}/|\varepsilon|=1.275$, where the maximum quality factor $\Omega T_{2}\approx 6$ is achieved. The spin is thermally initialized at $\langle\mathbf{S}\rangle_{\text{thermal}}=(-0.033,0,-0.188)$ and experiences a torque due to the time dependent exchange field. This torque causes the spin to precess in a plane with the normal vector $\mathbf{n}=(-\sin(4\pi/9),0,\cos(4\pi/9))$. b) Time evolution of the current showing up to six Rabi oscillation cycles before decaying to the steady state in the rotating frame. c) ESR signal, DC contribution, featuring a peak with a subtle Fano profile. The resonance frequency is slightly tuned with respect to the bare one to be $\Delta f+f_{0}=17.146$ GHz due to the energy dressing.
Figure 6: ESR in the Coulomb Blockade regime; in plane magnetic field. a) Bloch sphere representation of the spin components in the rotating frame for a magnetic field aligned with the plane ($\theta=\pi/2$) and a bias voltage of $eV_{\mathrm{DC}}/|\varepsilon|=1.275$, where the maximum quality factor $\Omega T_{2}\approx 36$ is achieved. The spin is thermally initialized at $\langle\mathbf{S}\rangle_{\text{thermal}}=(-0.19,0,0)$ and undergoes torque to the exchange field, leading to precession in a plane with the normal vector $\mathbf{n}=(0,0,1)$. b) Time evolution of the current, showing approximately 36 Rabi oscillations before decaying to the steady state in the rotating frame, closely matching the z-component of the spin, not the $x$-component. c) ESR signal, DC contribution, displaying a dip without Fano profile in the frequency domain. Since in-plane initialization leads to not dressing, the resonance frequency is $16.8$ GHz.
@@
The current in the rotating frame, as shown in Fig. 5b, directly reflects the Rabi oscillations of the spin, as it corresponds to the $z$-component of the spin. Notably, the DC contribution here is significantly smaller than that in Fig. 3b, as we are in the CB regime. This reduced DC contribution allows for a relatively long coherence time $T_{2}$ of approximately 35 ns without diminishing the contribution of the exchange field. An interesting aspect is observed in the ESR signal (Fig. 5c), which now exhibits a sharp peak instead of the dips seen earlier. This occurs because the ESR signal depends not only on the coherences but also on the populations of the spin states when the spin Rabi oscillates. On resonance, the populations tend toward a 50/50 distribution if the driving force is sufficiently strong, as expected in classical ESR simulations. At positive bias and spin polarization, the off-resonance population in the long-time limit leads to $\langle S_{z}\rangle>0$, product of, albeit small, spin accumulation, and consequently, the ESR signal on resonance manifests as a peak, see Eq. 12. Conversely, at negative bias, the signal flips to a dip, with an inverted Fano profile.
@@
We now turn to the time evolution of an in-plane initialized spin, Fig. 6a. The Rabi rate $\Omega_{\mathrm{FLT}}$ is maximized at $\theta=\pi/2$ radians, where the projection of the exchange field perpendicular to the spin is greatest, effectively maximizing the torque exerted by the exchange field. A notable feature in this simulation is the spin being pulled toward the positive $z$-axis of the Bloch sphere. This phenomenon can be attributed to the a non-negligible spin accumulation, which is expected given the DC bias’s proximity to the threshold $\varepsilon+U$. This residual spin accumulation explains a non-zero ESR signal even in the perpendicular configuration of the tip and spin—behavior, where typically no signal is found due to the homodyne detection [34].
@@
Since the current measures the $z$-component of the spin, it does not reflect the Rabi oscillations as clearly as in the out-of-plane case, as confirmed in Fig. 6b. However, the current’s maxima and minima in the rotating frame still correspond to the spin being in the ground and excited states, respectively, demonstrating controllability of the spin. It is worth noting that while our theory lacks a fully developed cotunneling framework, it captures key aspects of the detection mechanism, particularly homodyne detection. This is evident from the intensity minimum of the ESR signal in Fig. 6c, despite the higher quality factor $\Omega_{\mathrm{FLT}}T_{2}$ in the in-plane configuration compared to the out-of-plane case Fig. 5c. Thus, we conclude that while our model has limitations, it successfully incorporates the homodyne detection, and it aligns with experimental results, indicating that out-of-plane magnetic fields yield stronger ESR signals than in-plane configurations.

### v2

@@ III.3 Energy dressing
Returning to the exchange field, we observed that the FLT Rabi rate can be interpreted as the projection of the exchange field onto the axis perpendicular to the spin in the plane containing both the exchange field and the spin itself. However, we initially ignored the projection of this field along the eigenbasis axis. On it, the Hamiltonian is diagonal, and the projection of the exchange on this axis will change the energy of the states, usually referred to as dressing or Lamb shift. This projection provides a Zeeman-like splitting even in the absence of an external magnetic field, as long as the exchange field and spin are not perpendicular. When an external magnetic field is applied, the energy shift will modify the resonance frequency of the spectrum, shifting it relative to the bare resonance frequency given by $f_{0}=\Delta_{eg}/h=g\mu_{B}B/h$. Following this geometric interpretation, we define the energy (or frequency) shift as:
$\delta E(t)$ $=$ $h\delta f(t)=h(f_{0}^{\prime}-f_{0})=g\mu_{B}B_{\mathrm{exch}}(t)\cos\theta$
@@
where $f_{0}^{\prime}$ is the dressed resonance frequency. The time-dependent component in Eq. (LABEL:energy_dress) does not affect the spin dynamics, as it aligns with the eigenbasis axis, though we include it here for completeness.
@@
This relation was used to derive Eqs. (17a) and (17b). Equation (31) provides a straightforward method to effectively measure the Rabi frequency, given the knowledge of the frequency shift and the angle between the tip and the external magnetic field. Another important detail about the energy dressing is how it affects zero-harmonic and higher-harmonic signals. Writing Eq. (5) in Floquet components, as in [37, 38], shows that in general the resonance frequency per harmonic is $hf_{n\omega}=(\Delta_{eg}+h\delta f)/n$, which makes the energy shift dependent on the harmonic $n$ (or Floquet number). Interestingly, if we move to the perspective of the magnetic field being swept at a fixed frequency, the resonance magnetic field per harmonic will be $B_{n\omega}=h(nf-\delta f)/g\mu_{B}$. This implies that the shifting in the magnetic field domain is the same per harmonic since $\delta f$ is not altered by $n$. Following the equations above, the zero-harmonic feature in the frequency domain appears at zero frequency, while in the magnetic field domain, it does at $h\delta f/g\mu_{B}=-B_{\mathrm{exch}}\cos\theta$. As a consequence, this offset can be used to measure the exchange field indirectly, a proposal in line with other studies [46].
@@ IV Results
[…]
We have increased the intrinsic width of the impurity states to $\hbar/\tau_{c}=0.25\$meV, compared to the one used in [38], allowing exploration of the CB regime while maintaining $\hbar/\tau_{c}\ll|\varepsilon|,\ \varepsilon+U$. The driving amplitude is set to $A_{L}=50\%$, applied only to the polarized lead, consistent with the experimental setup. Finally, the Zeeman splitting provides a bare resonance frequency of approximately $f_{0}=16.8$ GHz, with an isotropic $g$-factor of 2 and $B=0.6$ T. The Zeeman energy that appears in the rates is not neglected in the numerical simulations, even though its contribution is extremely small.. The energy dressing will modify the resonance position to a value $f_{0}^{\prime}=\delta f+f_{0}$, see Eq. (LABEL:energy_dress), which is taken into account to perform on resonance simulations. Additional parameters are specified in figure captions.
IV.1 Spin accumulation, current and populations vs. DC bias under no AC voltage
@@
The effect of spin accumulation persists even in the CB regime, albeit weaker, vanishing completely at zero voltage. There is, however, a competition between the relaxation time, which ranges from 0.2 to 50 ns (Fig. 7a), and $\hbar/\langle\dot{s}_{z}\rangle_{\mathrm{acc}}$, leading to different results depending on the DC bias. In the CB regime, $\hbar/\langle\dot{s}_{z}\rangle_{\mathrm{acc}}\gg\tau_{\mathrm{rel}}$, so the spin evolution is dominated by the spin relaxation time, as the accumulation term acts extremely slowly on the QI. This opens the possibility of an FLT-ESR signal since the time dependent exchange field will have “time” to torque the spin coherently, producing Rabi oscillations. On the other hand, above the energy thresholds outside of the CB regime, both the relaxation and spin accumulation times are of the same order of magnitude, forcing the QI to be spin-accumulated in the eigenbasis axis before any coherent evolution occurs. Therefore, AC driving that leads to a time-dependent spin accumulation will provide an incoherent STT-ESR signal, as we will see in the following sections.
As a final remark, Fig. 2c shows the populations as a function of the DC bias, revealing a QI that is singly occupied throughout the CB regime; the empty and doubly occupied states become populated once the DC bias exceeds the thresholds at $\varepsilon$ and $\varepsilon+U$, respectively. This occurs when enough energy is available to remove or add an electron to the QI. The small overshoot in the ground-state population at negative bias (also present, but smaller, at positive bias) results from the positive spin polarization together with the excited state leaving the CB regime slightly before the ground state owing to the small Zeeman splitting. Consequently, the left lead, which is positively polarized, reduces the population of the up-like (excited) state, an effect difficult to visualize since this population already begins to decrease as $eV_{\mathrm{DC}}\to\varepsilon$, producing a sudden increase in the ground-state population (down-like) at $eV_{\mathrm{DC}}=\varepsilon+g\mu_{B}B/2$. At the positive threshold energy, the overshooting is less pronounced due to the inversion of populations at the $\varepsilon+U$ level. If the spin polarization is negative, the behavior of Fig. 2c is essentially reversed with respect to the DC bias, being the strongest sudden overshooting at $eV_{\mathrm{DC}}=\varepsilon+U-g\mu_{B}B/2$. Figure 2c also displays the absolute value of the current, Eq. (12), which remains constant outside the CB regime and decays rapidly within it, in agreement with the analytical expression in Appendix C, Eq. (36). Due to numerical limitations, the current at $V_{\mathrm{DC}}=0$ is quite small but non-zero.
@@ IV.2 ESR in the STT-driven regime
Figure 3: ESR outside of the CB regime; out of plane magnetic field. a) Bloch sphere representation of the spin components, on resonance with the AC field, in the lab frame for $\theta=\pi/18$ and positive $eV_{\mathrm{DC}}/|\varepsilon|=2$. The spin is initialized thermally at $\langle\mathbf{s}\rangle_{\mathrm{thermal}}/{{\color[rgb]{0,0,0}{\hbar}}\color[rgb]{0,0,0}}=(-0.033,0,-0.188)$ (red arrow). The positive DC bias causes the spin to flip to its spin polarization in the steady state as a consequence of the spin accumulation term. At the same time, the spin continuously performs Larmor oscillations even in the steady state. The Bloch sphere axes are reduced for clarity. b) Time-evolution of the current, Eq. (12), from the thermal state, reaching the steady state after 2 ns. c) Corresponding ESR electric current signal, $\Delta I=I_{L}-I_{\mathrm{BG}}$, in the STT-driven regime in the frequency domain. The resonance frequency is slightly different to the bare one at 16.8 GHz, being now $f_{0}^{\prime}=\Delta f+f_{0}=17.07$ GHz due to the energy dressing.
@@
Fig. 3b illustrates the spin dynamics, characterized by fast Larmor oscillations that persist even in the long-time limit. These oscillations occur in a plane perpendicular to the eigenbasis axis, parallel to the external magnetic field, see Fig. 3a, and arise from the time-dependent spin accumulation. This accumulation, proportional to the polarized current, attempts to torque the accumulated spin toward the spin polarization axis, which in result makes the crossed product $\mathbf{s}\times\mathbf{B}$ non-zero, forcing the spin to torque with an intensity proportional to $\sin\theta$. In this regime, the populations remain practically frozen, and the Larmor oscillations only modify the coherences, making the ESR current fully governed by them [38].
@@
Fig. 3c shows the corresponding frequency behavior. Although the ESR signal is significant, the background current is three orders of magnitude higher, resulting in an extremely low ESR signal relative to the background. This is expected in the STT-driven regime for the parameters used, as reflected in the low-quality factor $\Omega_{\mathrm{STT}}T_{2}$. The signal exhibits an asymmetric Fano profile due to the exchange field and the residual presence of homodyne detection. However, it becomes fully symmetric when $eV_{\mathrm{DC}}\gg\varepsilon+U$, where the exchange field vanishes (see Eq. (25)).
@@
For the in-plane configuration, Fig. 4a, the strongest Larmor oscillations and ESR signals (Figs. 4b and 4c)) are achieved. The Larmor precession now occurs in the $zy$-plane, again perpendicular to the magnetic field and the eigenbasis axis. If the STT-driven regime would follow the homodyne detection previously reported [42, 22], the ESR signal would be minimal or zero since the impurity spin and spin polarization od the tip are perpendicular, and the ESR signal is proportional to their scalar product. However, in this regime, this limitation does not hold as the spin will be driven towards the $z$ axis by the spin accumulation. After this initial push, the spin feels the torque from the external magnetic field, from the cross product $\langle\mathbf{s}\rangle\times\mathbf{B}$. This torque is maximized for in plane configurations, producing a stronger ESR signal compared to the out of plane situation. This interaction allows for the initial oscillatory evolution of the spin, observed in both Fig. 4a and Fig. 3a. Shortly thereafter, the spin settles into oscillations within the $zy$ plane in Fig. 4a, driven by the time-dependent spin accumulation and the torque exerted by the external magnetic field.
@@
The role of the exchange field in this scenario is analogous to the out-of-plane case; however, the ESR signal in Fig. 4c is nearly symmetrical. This symmetry arises because, when the spin is initialized along a Cartesian axis, $\langle s_{z}\rangle$ in Eq.(13) depends solely on either the populations or the coherences, but not on a combination of both. Consequently, no Fano profile emerges. Specifically, for an in-plane magnetic field, the $z$ component of the spin depends only on the real part of the coherences, while the $x$ component reflects the populations. Another way to view this is that the FLT-driven ESR is not detected/observed at $90$ degrees because of the Homodyne detection, and the measured response is due to the STT-driven ESR alone, which provides a fully symmetric ESR.
Figure 4: ESR outside of the CB regime; in-plane magnetic field. a) Bloch sphere representation of the spin components, on resonance with the AC field, in the lab frame for $\theta=\pi/2$ and $eV_{\mathrm{DC}}/|\varepsilon|=2$. The spin is thermally initialized at $\langle\mathbf{s}\rangle_{\mathrm{thermal}}/{{\color[rgb]{0,0,0}{\hbar}}\color[rgb]{0,0,0}}=(-0.19,0,0)$ (red arrow). Here, the spin accumulation forces alignment with the eigenbasis axis of the spin in the steady state, leading to continuous Larmor oscillations in the zyzy-plane. b) Time-evolution of the current, Eq. (12), from the thermal situation. c) ESR electric current signal in the STT regime for the case of a) in the frequency domain, $\Delta I=I_{L}-I_{\mathrm{BG}}$. Since in-plane initialization leads to not dressing, the resonance occurs at the bare one $f_{0}=16.8$ GHz.
Importantly, in the lab frame, with an initial in-plane configuration, the spin cannot polarize because the spin accumulation term is fully perpendicular to the external magnetic field. Note that in the in-plane configuration, the populations, measured in the $x$ axis, also remain unperturbed in the long-time limit as observed in the out-of-plane situation Fig. 3. Interestingly, in the rotating frame, the torque exerted by the spin accumulation becomes increasingly influential as the angle changes from the out-of-plane to the in-plane configuration, with the latter causing the spin to align with the $z$-axis in the rotating frame.
@@
The reader may notice that the ESR signal in both cases (Figs. 3c and 4c) are dips (a reduction of the current on-resonance) which is a direct result of the STT-driven regime at positive bias. In this regime, the coherences peak on resonance while remaining zero off resonance. According to Eq. (12), this behavior produces a dip in the signal. Conversely, under a negative DC bias ($eV_{\mathrm{DC}}<\varepsilon$), the current becomes negative and the ESR signal will flip from a dip to a peak. However, the ESR signal cannot flip by merely changing the sign of the spin polarization because an additional sign change occurs in the $z$ component of the spin, effectively preserving the overall response. Higher order tunneling processes could induce additional flipping of the signal though.
@@ IV.3 ESR in the FLT-driven regime
We now shift our focus to the CB regime, where FLT-driven resonance dominates. This regime favors quantum-coherent manipulation of the spin over the spin polarization of the impurity spin since the spin accumulation term is significantly minimized. This reduces the impact of the completely decoherent polarized current on the impurity, enabling the spin to be used as a qubit. FLT driven regime is basically equivalent to conventional ESR, where a time dependent magnetic field induces a resonant transition. As highlighted in the theoretical discussion, this allows for quality factors $\Omega_{\mathrm{FLT}}T_{2}$ to reach their upper limits, constrained only by the ability to enhance $T_{2}$ for a fixed Rabi rate. Notably, this enhancement is strongly dependent on the DC bias, making it the critical factor in achieving high coherence for spin manipulation.
Figure 5: ESR in the Coulomb Blockade regime; out of plane magnetic field. a) Bloch sphere representation of the spin components in the rotating frame for a magnetic field tilted by $\theta=\pi/18$ and a bias voltage of $eV_{\mathrm{DC}}/|\varepsilon|=1.275$, where the maximum quality factor $\Omega T_{2}\approx 6$ is achieved. The spin is thermally initialized at $\langle\mathbf{s}\rangle_{\text{thermal}}=(-0.033,0,-0.188)$ and experiences a torque due to the time dependent exchange field. This torque causes the spin to precess in a plane with the normal vector $\mathbf{n}=(-\sin(4\pi/9),0,\cos(4\pi/9))$. b) Time evolution of the current, Eq. (12), showing up to six Rabi oscillation cycles before decaying to the steady state in the rotating frame. c) ESR signal, $\Delta I=I_{L}-I_{\mathrm{BG}}$, DC contribution, featuring a peak with a subtle Fano profile. The resonance frequency is slightly tuned with respect to the bare one to be $f_{0}^{\prime}=\Delta f+f_{0}=17.146$ GHz due to the energy dressing.
@@
The current in the rotating frame, as shown in Fig. 5b, directly reflects the Rabi oscillations of the spin, as it corresponds to the $z$-component of the spin. Notably, the DC contribution here is significantly smaller than that in Fig. 3b, as we are in the CB regime. This reduced DC contribution allows for a relatively long coherence time $T_{2}$ of approximately 35 ns without diminishing the contribution of the exchange field. An interesting aspect is observed in the ESR signal (Fig. 5c), which now exhibits a sharp peak instead of the dips seen earlier. This occurs because the ESR signal depends not only on the coherences but also on the populations of the spin states when the spin Rabi oscillates. On resonance, the populations tend toward a 50/50 distribution if the driving force is sufficiently strong, as expected in classical ESR simulations. At positive bias and spin polarization, the off-resonance population in the long-time limit leads to $\langle s_{z}\rangle>0$, a product of, albeit small, spin accumulation, and consequently, the ESR signal on resonance manifests as a peak, see Eq. (12). Conversely, at negative bias, the signal flips to a dip, with an inverted Fano profile.
@@
We now turn to the time evolution of an in-plane initialized spin, Fig. 6a. The Rabi rate $\Omega_{\mathrm{FLT}}$ is maximized at $\theta=\pi/2$ radians, where the projection of the exchange field perpendicular to the spin is greatest, effectively maximizing the torque exerted by the exchange field. A notable feature in this simulation is the spin being pulled toward the positive $z$-axis of the Bloch sphere. This phenomenon is attributed to a non-negligible spin accumulation, which is expected given the DC bias’s proximity to the threshold $\varepsilon+U$. This residual spin accumulation explains a non-zero ESR signal even in the perpendicular configuration between the tip’s magnetization and the impurity spin, where typically no signal is found due to the homodyne detection rendering the FLT-driven component unobservable, even though the Rabi rate is maximized [42, 22]. Since this measured response is due to the remanent contribution from the STT-driven regime alone, a fully symmetric ESR is obtained, see Fig. 6c, although without the strong damping effect of the STT regime since $\hbar/\langle\dot{s}_{z}\rangle_{\mathrm{acc}}\gg T_{2}$.
Figure 6: ESR in the Coulomb Blockade regime; in plane magnetic field. a) Bloch sphere representation of the spin components in the rotating frame for a magnetic field aligned with the plane ($\theta=\pi/2$) and a bias voltage of $eV_{\mathrm{DC}}/|\varepsilon|=1.275$, where the maximum quality factor $\Omega T_{2}\approx 36$ is achieved. The spin is thermally initialized at $\langle\mathbf{s}\rangle_{\text{thermal}}=(-0.19,0,0)$ and undergoes torque to the exchange field, leading to precession in a plane with the normal vector $\mathbf{n}=(0,0,1)$. b) Time evolution of the current, Eq. (12), showing approximately 36 Rabi oscillations before decaying to the steady state in the rotating frame, closely matching the z-component of the spin, not the $x$-component. c) ESR signal, $\Delta I=I_{L}-I_{\mathrm{BG}}$, DC contribution, displaying a dip without a Fano profile in the frequency domain. Since in-plane initialization leads to not dressing, the resonance frequency is $16.8$ GHz.
@@
Since the current measures the $z$-component of the spin, it does not reflect the Rabi oscillations as clearly as in the out-of-plane case, as confirmed in Fig. 6b. However, the current’s maxima and minima in the rotating frame still correspond to the spin being in the ground and excited states, respectively, demonstrating controllability of the spin, visible only due to the small driving contribution from STT. Further reducing this contribution by lowering the DC bias would completely suppress the ESR signal. It is worth noting that while our theory lacks a fully developed cotunneling framework, it captures key aspects of the detection mechanism, particularly homodyne detection. This is evident from the intensity minimum of the ESR signal in Fig. 6c, despite the higher quality factor $\Omega_{\mathrm{FLT}}T_{2}$ in the in-plane configuration compared to the out-of-plane case Fig. 5c. Thus, we conclude that while our model has limitations, it successfully incorporates the homodyne detection, and it aligns with experimental results, indicating that out-of-plane magnetic fields yield stronger ESR signals than in-plane configurations.